OpenStudy (anonymous):

Find the area of a trapezoid with bases of 18 cm and 26 cm and a height of 11 cm Find the area of a trapezoid with bases of 18 cm and 26 cm and a height of 11 cm @Mathematics

OpenStudy (amistre64):

area of traps is the average of the bases times the height

OpenStudy (amistre64):

18+26 = 44 44/2 = 22 22*11 = 242

OpenStudy (amistre64):

;) yeah, that last bit is useful on tests that like to trick yah